# scE2TM improves single-cell embedding interpretability and reveals cellular perturbation signatures

### v2

@@ 1 Introduction
We present scE2TM (single-cell External knowledge-guided Embedding clustering regularization Topic Model), a deep generative model for large-scale single-cell transcriptomic analysis (Fig. 1). scE2TM introduces two key innovations to overcome interpretation collapse. First, Embedding Clustering Regularization (ECR) [wu2023effective] encourages topic diversity by regularizing topic embeddings as clustering centers and gene embeddings into clustering samples, with cluster soft assignments modeled through Optimal Transport (OT) and enhances representational learning of distinct biological processes. Second, scE2TM integrates external knowledge from single-cell foundation models through a Cross-view Encoder (CVE), shifting from internal guidance paradigms to scalable external knowledge integration. To systematically evaluate scE2TM, we employed ten quantitative interpretability metrics that, combined with clustering metrics, enable rigorous assessment of topic model performance. Benchmarking across 20 datasets revealed that scE2TM mitigates interpretation collapse while existing methods suffer. Critically, clustering performance and interpretability are mostly uncorrelated–demonstrating that previous models sacrifice bioloical insights while optimizing clustering accuracy. scE2TM achieved state-of-the-art performance in both aspects. We then showcase real-world applications of scE2TM in three case studies. In a human pancreas scRNA-seq dataset, scE2TM captured insulin biosynthesis pathways and programs reflecting key cellular identities. Topic-based perturbations in interferon (IFN)-stimulated PBMCs successfully transformed cellular states, validating capacity to simulate cellular responses to exogenous stimuli. In a melanoma scRNA-seq dataset, scE2TM-derived topics generalized to TCGA cohorts with topic activity significantly correlating with patient survival, demonstrating clinical relevance. Collectively, these results establish scE2TM’s as a framework for extracting mechanistically insightful and clinically actionable patterns from single-cell transcriptomics.
@@ 2.4 Quantitative benchmark of model interpretability
We investigated the top ten genes for each scE2TM topic and calculated their frequency of occurrence across topics. Our results show the ten most frequent genes, including INS, CEACAM7, and UCA1 (Supplementary Table LABEL:supt3), all experimentally validated or reported in the literature as significantly associated with pancreatic function or related disease mechanisms [andrali2008glucose, raj2021ceacam7, chen2016long]. Pathway enrichment analysis identifies many enriched pathways closely related to pancreatic function for each topic (Benjamin-Hochberg (BH) false discovery rate 0.05; Supplementary Fig. LABEL:supf7). Enrichment of “REACTOME_INSULIN_PROCESSING” and “VANGURP_PANCREATIC_BETA_CELL” highlights the model’s ability to capture key pancreatic functions, including insulin biosynthesis and cell-specific molecular characteristics (Fig. 5c,d). Cosine similarity analysis quantifies the geometric structure of the embedding space: genes with higher topic weights display greater embedding similarity to their corresponding topics (Fig. 5e). The analysis of human pancreas data establishes that scE2TM recovers biologically coherent topic-gene structures in the embedding space and successfully identifies key pancreatic functional genes and pathways, validating its capability to uncover mechanistically relevant biology from single-cell transcriptomes.
@@ 2.6 Topic-based perturbation reveals distinct interferon response in cell-state transitions
We applied scE2TM to the PBMC dataset from Kang et al [kang2018multiplexed], which includes single-cell RNA-seq data from lupus patient PBMCs treated with IFN-, a type I IFN (Section 4.6). scE2TM effectively separates known cell types and treatment conditions while identifying several topics specifically associated with IFN stimulation (Fig. 6a,b). Differential expressed topic analysis identifies topics 59 and 76 (IFN-specific topic) as significantly up-regulated in IFN-stimulated cells versus control cells (BH-adjusted -value 0.05; Fig. 6c; Supplementary Fig. LABEL:supf8). Topic 76 exhibits the highest activation levels in FCGR3A+Mono cells. The top genes of topic 76 are significantly enriched for for both FCGR3A+Mono-specific up-regulate genes (Supplementary Fig. LABEL:supf9a) and type I IFN response pathway components (Supplementary Fig. LABEL:supf9b).
@@ 2.7 Topic-based perturbation shifts melanoma cells towards normal state
To assess whether the perturbed malignant cells shift towards the normal state, we trained an SVM classifier to predict cell states (malignant or normal) (Section 4.9). As perturbation intensity increases by progressive suppression of malignant-specific topics, a greater proportion of malignant cells become classified as normal (Fig. 8b), indicating that these topics are critical for maintaining the malignant phenotype and might represent key drivers of malignant transformation. To visualize these transitions, we projected perturbed malignant cells into UMAP space. With increasing perturbation of the malignant-specific topics, the perturbed cells gradually converge toward the normal cell cluster (Fig. 8c), whereas perturbing non-malignant-specific topics topics results in cells remaining clustered among the original malignant cells (Supplementary Fig. LABEL:supf13). Furthermore, we found that perturbing the top- malignant-specific topics induced detectable shifts in cellular states, and that these shifts became more pronounced as the number of perturbed malignant-specific topics increased (Supplementary Fig. LABEL:supf14). This observation indicates that the malignant-specific topics identified by scE2TM represent key transcriptional axes of tumor progression, and perturbing them reverses malignant cells toward normal phenotypes.
To assess the clinical relevance of our analysis, we further analyzed melanoma patient data from The Cancer Genome Atlas (TCGA) [hoadley2018cell]. The up-regulated genes identified in the melanoma dataset show a significant negative correlation with survival rates in TCGA melanoma cohort patients (Fig. 8d; -value = 1.6e-4), confirming that perturbation of the target indeed impacts key signaling pathways in melanoma biology. We then applied the melanoma-trained scE2TM model to the TCGA melanoma gene expression data and examined the associations between the previously identified malignant-specific topics and patient outcomes. The expression of malignant-specific topics shows significant associations with survival (Fig. 8e,f), demonstrating that scE2TM-identified topics capture clinically relevant gene programs that generalize across independent patient cohorts. Together, scE2TM reveals malignancy-specific topics and can be robustly extended to independent TCGA melanoma studies. These findings suggest that interpretable topic characterization holds promise as a bridge linking single-cell transcriptomics to cancer clinical prognosis prediction.
@@ 3 Discussion
Second, to address interpretation collapse, we adopt embedding cluster regularization (ECR), which treats topics as clustering centroids and genes as samples. Through optimal transport, the ECR ensures that each topic serves as the anchors of a distinct gene cluster. Quantitative experiments show scE2TM’s substantial improvement over benchmark models, with increases of 8.6% in interpretation purity, 66.5% in topic quality, 27.5% in GSEA quality, and 16.1% in ORA quality. Qualitative experiments further confirm that scE2TM topics capture cell type-specific and functionally relevant biological signals. Our correlation analysis between clustering and interpretability metrics reveals that high clustering performance does not guarantee biologically meaningful interpretations, underscoring the importance of dedicated interpretability assessment.
@@
Future extensions of our study will focus on three directions. First, we will collect tissue- or cell type-specific pathway datasets to enable a more accurate evaluation of the model’s interpretability. Although scE2TM identifies topics enriched for diverse and coherent biological pathways, several interpretability metrics (e.g., ORAQ and GSEAQ) are evaluated against all known pathways and therefore cannot determine how well these pathways reflect the underlying cellular characteristics. Collecting biological pathway datasets with tissue- or cell type–specific information is a promising direction for improving our interpretability evaluation benchmarks. Nevertheless, we anticipate that diversity and consistency will remain valid assessment criteria. Secondly, existing single-cell embedding topic models generate global topic-gene dependencies for entire datasets, potentially overlooking cell-specific characteristics. Developing hierarchical topic structures that generate topics at multiple granularities would enable more refined interpretations focused on individual cell features [wu2024survey]. Finally, with the increasing prevalence of multimodal single-cell assays, extending scE2TM and our interpretability metrics to multimodal data would enable more comprehensive biological signal capture across modalities [ZHOU2023100563, coleman2025resolving, cui2025towards].

### v3

@@ 1 Introduction
We present scE2TM (single-cell External knowledge-guided Embedding clustering regularization Topic Model), a deep generative model for large-scale single-cell transcriptomic analysis (Fig. 1). scE2TM introduces two key innovations to overcome interpretation collapse. First, Embedding Clustering Regularization (ECR) [wu2023effective] encourages topic diversity by regularizing topic embeddings as clustering centers and gene embeddings into clustering samples, with cluster soft assignments modeled through Optimal Transport (OT) and enhances representational learning of distinct biological processes. Second, scE2TM integrates external knowledge from single-cell foundation models through a Cross-view Encoder (CVE), shifting from internal guidance paradigms to scalable external knowledge integration. To systematically evaluate scE2TM, we employed ten quantitative interpretability metrics that, combined with clustering metrics, enable rigorous assessment of topic model performance. Benchmarking across 20 datasets revealed that scE2TM mitigates interpretation collapse while existing methods suffer. Critically, clustering performance and interpretability are mostly uncorrelated—demonstrating that previous models sacrifice bioloical insights while optimizing clustering accuracy. scE2TM achieved state-of-the-art performance in both aspects. We then showcase real-world applications of scE2TM in three case studies. In a human pancreas scRNA-seq dataset, scE2TM captured insulin biosynthesis pathways and programs reflecting key cellular identities. Topic-based perturbations in interferon (IFN)-stimulated PBMCs successfully transformed cellular states, validating capacity to simulate cellular responses to exogenous stimuli. In a melanoma scRNA-seq dataset, scE2TM-derived topics generalized to TCGA cohorts with topic activity significantly correlating with patient survival, demonstrating clinical relevance. Collectively, these results establish scE2TM’s as a framework for extracting mechanistically insightful and clinically actionable patterns from single-cell transcriptomics.
@@ 2.4 Quantitative benchmark of model interpretability
We investigated the top ten genes for each scE2TM topic and calculated their frequency of occurrence across topics. Our results show the ten most frequent genes, including INS, CEACAM7, and UCA1 (Supplementary Table LABEL:supt3), all experimentally validated or reported in the literature as significantly associated with pancreatic function or related disease mechanisms [andrali2008glucose, raj2021ceacam7, chen2016long]. Pathway enrichment analysis identifies many enriched pathways closely related to pancreatic function for each topic (Benjamin-Hochberg (BH) false discovery rate 0.01; Supplementary Fig. LABEL:supf7). Enrichment of “REACTOME_INSULIN_PROCESSING” and “VANGURP_PANCREATIC_BETA_CELL” highlights the model’s ability to capture key pancreatic functions, including insulin biosynthesis and cell-specific molecular characteristics (Fig. 5c,d). Cosine similarity analysis quantifies the geometric structure of the embedding space: genes with higher topic weights display greater embedding similarity to their corresponding topics (Fig. 5e). The analysis of human pancreas data establishes that scE2TM recovers biologically coherent topic-gene structures in the embedding space and successfully identifies key pancreatic functional genes and pathways, validating its capability to uncover mechanistically relevant biology from single-cell transcriptomes.
@@ 2.6 Topic-based perturbation reveals distinct interferon response in cell-state transitions
We applied scE2TM to the PBMC dataset from Kang et al [kang2018multiplexed], which includes single-cell RNA-seq data from lupus patient PBMCs treated with IFN-, a type I IFN. scE2TM effectively separates known cell types and treatment conditions while identifying several topics specifically associated with IFN stimulation (Fig. 6a,b). Differential expressed topic analysis (Section 4.6) identifies topics 59 and 76 (IFN-specific topic) as significantly up-regulated in IFN-stimulated cells versus control cells (BH-adjusted -value 0.01; Fig. 6c; Supplementary Fig. LABEL:supf8). Topic 76 exhibits the highest activation levels in FCGR3A+Mono cells. The top genes of topic 76 are significantly enriched for for both FCGR3A+Mono-specific up-regulate genes (Supplementary Fig. LABEL:supf9a) and type I IFN response pathway components (Supplementary Fig. LABEL:supf9b).
@@ 2.7 Topic-based perturbation shifts melanoma cells towards normal state
To assess whether the perturbed malignant cells shift towards the normal state, we trained an SVM classifier to predict cell states (malignant or normal) (Section 4.9). As perturbation intensity increases by progressive suppression of malignant-specific topics, a greater proportion of malignant cells become classified as normal (Fig. 8b), indicating that these topics are critical for maintaining the malignant phenotype and might represent key drivers of malignant transformation. To visualize these transitions, we projected perturbed malignant cells into UMAP space. With increasing perturbation of the malignant-specific topics, the perturbed cells gradually converge toward the normal cell cluster (Fig. 8c), whereas perturbing non-malignant-specific topics results in cells remaining clustered among the original malignant cells (Supplementary Fig. LABEL:supf13). Furthermore, we found that perturbing the top- malignant-specific topics induced detectable shifts in cellular states, and that these shifts became more pronounced as the number of perturbed malignant-specific topics increased (Supplementary Fig. LABEL:supf14). This observation indicates that the malignant-specific topics identified by scE2TM represent key transcriptional axes of tumor progression, and perturbing them reverses malignant cells toward normal phenotypes.
To assess the clinical relevance of our analysis, we further analyzed melanoma patient data from The Cancer Genome Atlas (TCGA) [hoadley2018cell]. The up-regulated genes identified in the melanoma dataset show a significant negative correlation with survival rates in TCGA melanoma cohort patients (Fig. 8d; -value = 1.6e-4), confirming that perturbation of the target indeed impacts key signaling pathways in melanoma biology. We then applied the melanoma-trained scE2TM model to the TCGA melanoma gene expression data and examined the associations between the previously identified malignant-specific topics and patient outcomes. The expression of malignant-specific topics shows significant associations with survival (Fig. 8e,f), demonstrating that scE2TM-identified topics capture clinically relevant gene programs that generalize across independent patient cohorts. Together, scE2TM reveals malignant-specific topics and can be robustly extended to independent TCGA melanoma studies. These findings suggest that interpretable topic characterization holds promise as a bridge linking single-cell transcriptomics to cancer clinical prognosis prediction.
@@ 3 Discussion
Second, to address interpretation collapse, we adopt ECR, which treats topics as clustering centroids and genes as samples. Through optimal transport, the ECR ensures that each topic serves as the anchors of a distinct gene cluster. Quantitative experiments show scE2TM’s substantial improvement over benchmark models, with increases of 8.6% in interpretation purity, 66.5% in topic quality, 27.5% in GSEA quality, and 16.1% in ORA quality. Qualitative experiments further confirm that scE2TM topics capture cell type-specific and functionally relevant biological signals. Our correlation analysis between clustering and interpretability metrics reveals that high clustering performance does not guarantee biologically meaningful interpretations, underscoring the importance of dedicated interpretability assessment.
@@
Future extensions of our study will focus on three directions. First, we will collect tissue- or cell type-specific pathway datasets to enable a more accurate evaluation of the model’s interpretability. Although scE2TM identifies topics enriched for diverse and coherent biological pathways, several interpretability metrics (e.g., ORAQ and GSEAQ) are evaluated against all known pathways and therefore cannot determine how well these pathways reflect the underlying cellular characteristics. Collecting biological pathway datasets with tissue- or cell type-specific information is a promising direction for improving our interpretability evaluation benchmarks. Nevertheless, we anticipate that diversity and consistency will remain valid assessment criteria. Secondly, existing single-cell embedding topic models generate global topic-gene dependencies for entire datasets, potentially overlooking cell-specific characteristics. Developing hierarchical topic structures that generate topics at multiple granularities would enable more refined interpretations focused on individual cell features [wu2024survey]. Finally, with the increasing prevalence of multimodal single-cell assays, extending scE2TM and our interpretability metrics to multimodal data would enable more comprehensive biological signal capture across modalities [ZHOU2023100563, coleman2025resolving, cui2025towards].